OpenStudy (anonymous):

integrate 3sqrt2 to 6 1/(x^3sqrt(x^2-9))dx

OpenStudy (anonymous):

try \(=x^2\) and \(du=2xdx\) as a start

OpenStudy (anonymous):

should i not use the trig sub x=3secø?

OpenStudy (anonymous):

oh damn sorry nmv i didn't see it was a fraction tilt

OpenStudy (anonymous):

yes, you should i make a mistake

OpenStudy (anonymous):

where do i go after i sub to (3secøtanødø)/(3secøtan^2ø)

OpenStudy (anonymous):

\[\frac{ 3\sec \theta \tan \theta d \theta \ }{ 3\sec \theta \tan^2 \theta }\]

OpenStudy (anonymous):

problem started as \[\int\limits_{3\sqrt{2}}^{6} \frac{ 1 }{ x^3\sqrt{x^2-9} }dx\] from which i subbed \[x=3\sec \theta\] \[dx = 3(\sec \theta \tan \theta) d \theta\] to get to \[\frac{ 3\sec \theta \tan \theta d \theta }{ 3\sec^3 \theta \tan^2 \theta}\]

OpenStudy (zarkon):

try again

OpenStudy (anonymous):

27 sec ^3 theta *

OpenStudy (zarkon):

\[x^3\sqrt{x^2-9}\] \[=(3\sec(\theta))^3\sqrt{(3\sec(\theta))^2-9}\] \[=27\sec^3(\theta)\sqrt{9(\sec^2(\theta)-1)}\] \[=81\sec^3(\theta)\sqrt{\tan^2(\theta)}\] \[=81\sec^3(\theta)\tan(\theta)\]

OpenStudy (anonymous):

ok thats right I see where I messed up but then where do I go with the problem from there?

OpenStudy (zarkon):

simplify

OpenStudy (zarkon):

then integrate

OpenStudy (anonymous):

i can take the 3/81 out which leaves just \[du=\sec \theta \tan \theta\] in the numerator if my \[u=\sec \theta\] and that means my denom is \[u^2du\] then there are 2 du or an extra tangent?

OpenStudy (zarkon):

\[\frac{3\sec(\theta)\tan(\theta)}{81\sec^3(\theta)\tan(\theta)}d\theta\] \[=\frac{1}{27\sec^2(\theta)}d\theta=\frac{1}{2}\cos^2(\theta)d\theta\]

OpenStudy (zarkon):

\[=\frac{1}{27\sec^2(\theta)}d\theta=\frac{1}{27}\cos^2(\theta)d\theta\]

OpenStudy (anonymous):

wow, i totally got caught up in the trig to not see that reduce... thank you very much for your patience zarkon

OpenStudy (zarkon):

you know what to do from here?

OpenStudy (anonymous):

yeah